(x+h)(x+h)(x+h)= ?
What's the method for solving this equation? I know FOIL for (x+h)(x+h)= , but not the method for cubing.

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

(x+h)(x+h)(x+h)= ?
= (x^2+2hx+h^2)(x+h)
= (x^2+2hx+h^2)x + (x^2+2hx+h^2)h
= x^3+2hx^2+h^2x + hx^2+2h^2x + h^3
= x^3 + (2x+h)x^2 + (h^2+2h^2)x + h^3
= x^3 + (3h)x^2 + (3h^2)x + h^3
==============
Cheers,
Stan H.

Answer by checkley71(8403)

FOIL works for 2 equations the old fashioned multiplacation works for any number of equations multiplied as follows.
x+h
x+h
----------------
x^2+hx+hx+h^2 combine like terms
x^2+2hx+h^2
x+h
---------------
x^3+2hx^2+h^2x+hx^2+2h^2x+h^3 combine like terms
x^3+3hx^2+3h^2x+h^3 answer